40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 18 20:17:17 2014, Jenne, Update, LSC, ALS noisy ALS_1kHzRMSnoise.pdfALS_oolNoise.pdf
    Reply  Mon Apr 21 22:53:16 2014, manasa, Update, LSC, ALS noise ALSX_FreeInLoop.jpgALSXY_inLoop.jpg
       Reply  Tue Apr 22 01:11:42 2014, Jenne, Update, LSC, TRY 60Hz noise TRY_60Hz_noise_21Apr2014.pdf
          Reply  Tue Apr 22 21:54:50 2014, manasa, Update, LSC, TRY 60Hz noise 60HzTRY.jpg
             Reply  Tue Apr 22 22:49:10 2014, rana, Update, LSC, TRY 60Hz noise 
                Reply  Wed Apr 23 19:58:00 2014, manasa, Update, LSC, TRY 60Hz noise 
                   Reply  Tue May 6 18:32:14 2014, steve, Update, LSC, TRY 60Hz noise hunt ETMY-ISCT_EISOL.jpg
                      Reply  Wed May 7 14:01:36 2014, steve, Update, LSC, TRY 60Hz noise hunt 
                         Reply  Thu May 8 17:25:17 2014, steve, Update, LSC, TRY 60Hz noise hunt 
                         Reply  Fri May 16 16:08:12 2014, steve, Update, LSC, Isolating base plates ISObaseplates.jpg
                Reply  Wed Apr 23 23:48:30 2014, manasa, Update, LSC, Y end whitening board IMG_1378.JPGIMG_1379.JPGIMG_1378.JPG
                   Reply  Thu Apr 24 14:23:09 2014, not manasa, Update, LSC, Y end whitening board 
                      Reply  Thu Apr 24 23:55:31 2014, Koji, Update, LSC, Y end whitening board P4245550.JPG
                         Reply  Fri Apr 25 10:43:57 2014, Koji, Update, LSC, (Fixed) Y end whitening board D990399-B_40m.pdfP4245552.JPGP4245553.JPGP4245551.JPG
                            Reply  Thu Dec 11 18:12:46 2014, ericq, Update, LSC, (Fixed) Y end whitening board 
                               Reply  Thu Dec 11 22:44:23 2014, rana, Update, LSC, QPD screens 
          Reply  Thu Apr 24 00:11:35 2014, Jenne, Update, LSC, Yend shutter back. 
       Reply  Sun Apr 27 20:26:19 2014, Koji, Update, LSC, Phase Tracker servo characterization ALSX_PTTF.pdfALSY_PTTF.pdf
          Reply  Sun Apr 27 21:30:59 2014, Koji, Update, LSC, ALS servo characterization ALSX_OLTF.pdfALSY_OLTF.pdf
             Reply  Mon Apr 28 10:24:10 2014, Koji, Update, LSC, error signal characterization ALSX_SPE.pdfALSY_SPE.pdf
                Reply  Mon Apr 28 10:34:51 2014, Koji, Update, LSC, new ALS servo design ALSX_OLTF_new2.pdfALSY_OLTF_new2.pdf
                   Reply  Mon Apr 28 10:48:48 2014, Koji, Update, LSC, new ALS servo design: comparison ALSX_OLTF_new.pdfALSY_OLTF_new.pdf
                      Reply  Mon Apr 28 10:59:54 2014, Koji, Update, LSC, New ALS servo design: expected error signals ALSX_SPE_new.pdfALSY_SPE_new.pdf
                         Reply  Mon Apr 28 11:08:11 2014, Koji, Update, LSC, New ALS servo design: expected error signals 140421_ALS_servo.zip
Message ID: 9921     Entry time: Wed May 7 14:01:36 2014     In reply to: 9918     Reply to this: 9933   9965
Author: steve 
Type: Update 
Category: LSC 
Subject: TRY 60Hz noise hunt 

Quote:

This is an effort to get rid of our ground loops  by isolating the electronic components from the optical table.

Aluminum mounting base plates of Thorlabs BA2 and Newport B-2 were replaced by plates or post made out of delrin material.

This is an insulator. DELRIN base plates were installed 6 places. The oplev-qpd has Nylon base plate.

The NPRO and HE/NE lasers are not isolated from the table. S8 and S9

I'm not sure about the doubling oven S10 

The optical table is grounded at G11  through  ~1 Mohms to the ETMY chamber.

Alignment touch up needed   at all D-marked component!

 

 Attachment appendix:

 

D: component delrin isolated

N: component nylon isolated ( or Delrin )

S: component shell is shorting to optical table (except oven)

G:  optical table ground

 

I failed to maximize TRY the pds.

ELOG V3.1.3-